# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v1

@@ 5 MUMFORD EXPERIMENTS
We first evaluated our method on the Mandl (1980) and Mumford (2013b) city datasets, two popular benchmarks for evaluating NDP algorithms (Mumford, 2013a; John et al., 2014; Kılıç and Gök, 2014; Ahmed et al., 2019). The Mandl dataset is one small synthetic city, while the Mumford dataset consists of four synthetic cities, labelled Mumford0 through Mumford3, and it gives values of $S$, $MIN$, and $MAX$ to use when benchmarking on each city. The values $n$, $S$, $MIN$, and $MAX$ for Mumford1, Mumford2, and Mumford3 are taken from three different real-world cities and their existing transit networks, giving the dataset a degree of realism. Details of these benchmarks are given in Table 1.
@@
[…]
@@ 5.1 Comparison with Baseline Evolutionary Algorithm
We run each algorithm under consideration on all five of these synthetic cities over eleven different values of $\alpha$, ranging from $0.0$ to $1.0$ in increments of $0.1$. This lets us observe how well the different methods perform under a range of possible preferences, from the extremes of the operator perspective ($\alpha=0.0$) and passenger perspective ($\alpha=1.0$) to a range of intermediate perspectives. The constraint weight $\beta=5.0$ in all experiments, the same value used in training the policies. We found that this $\beta$ was sufficient: none of the transit networks produced in our experiments violated any of the constraints in section 3.
We perform ten runs of each method with ten different random seeds. For algorithms that make use of a learned constructor policy, ten such policies were trained with the same set of random seeds (but using the same training dataset), and each was used in the run with the corresponding random seed. We report the mean and standard deviation of total route time $C_{o}$ and average trip time $C_{p}$ for each run.
We first compared our neural evolutionary algorithm (NEA) with our baseline evolutionary algorithm (EA). We also compare both with the initial solutions from LC-100, to see how much each of EA and NEA improve the solution. In the EA and NEA runs, the same parameter settings of $B=10,IT=400,E=10$ were used. Table 2 gives statistics of the final cost achieved by each algorithm on Mandl and the four Mumford cities for the operator perspective ($\alpha=0.0$), the passenger perspective ($\alpha=1.0$), and a balanced perspective ($\alpha=0.5$). We see that on Mandl, the smallest of the five cities, EA and NEA perform virtually identically for the operator and passenger perspectives, while NEA offers a slight improvement for the balanced perspective. But for the larger cities of the Mumford dataset, NEA performs considerably better than EA for both the operator and balanced perspectives.
For the passenger perspective NEA’s advantage over EA is smaller, but it still outperforms EA on Mumford1 and Mumford3 and performs comparably on Mumford2. On Mumford3, the largest of the five cities, NEA solutions EA have 14% lower average cost for the passenger perspective and 6% lower for the balanced perspective. For the operator perspective, the difference is much less significant: NEA’s solutions have 1% lower average cost, but the average costs of NEA and EA are within one standard deviation of each other.
@@
Figure 3 displays the values of average trip time $C_{p}$ and total route time $C_{o}$ achieved by each algorithm on the three largest Mumford cities, the three which are each based on a real city’s statistics, over the full range of $\alpha$ values. There is a necessary trade-off between these two values, and as we expect, as $\alpha$ increases, we observe that $C_{o}$ increases while $C_{p}$ decreases for each method. We also observe that for intermediate values of $\alpha$, NEA pushes $C_{o}$ much lower than LC-100, at the cost of increases in $C_{p}$. This behaviour is desirable, because as the figure shows, NEA achieves an overall larger range of outcomes than LC-100 - for any solution from LC-100, there is some value of $\alpha$ for which NEA produces a strictly dominant solution. Figure 3 also shows results for an additional algorithm, LC-40k, which will be discussed in subsection 5.2.
On all three cities, we observe a common pattern: EA and NEA perform very similarly at $\alpha=1.0$ (the leftmost point on each curve), but a significant performance gap forms as $\alpha$ decreases - consistent with what is shown by Table 2. We also observe that LC-100 performs better in terms of $C_{p}$ than $C_{o}$, with its points clustered higher and more leftwards than most of the points with corresponding $\alpha$ values on the other curves. Both EA and NEA make only very small decreases in $C_{p}$ on LC-100’s initial solutions at $\alpha=1.0$, and they do so by increasing $C_{o}$ considerably.
@@
Figure 3: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis), across values of $\alpha$ evenly spaced over the range $[0,1]$, from $\alpha=0.0$ at right to $1.0$ at left, for transit networks from LC-100, LC-40k, EA, and NEA. Points are the average over 10 random seeds, and bars around each point indicate one standard deviation. A line links every two points with consecutive $\alpha$ values. Lower values of $C_{o}$ and $C_{p}$ are better, so the lower-left direction in each plot represents improvement.
@@
Figure 4: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis) achieved by all-1 NEA, plotted along with EA and NEA for comparison.
@@
Figure 5: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis) achieved by PC-EA, plotted along with LC-100, EA, and NEA for comparison.
@@ 5.2 Ablation Studies
To better understand the contribution of various components of our method, we performed a set of three ablation studies, described below. These were conducted over the three realistic Mumford cities and over the same range of $\alpha$ values as in subsection 5.1.
@@ 5.2.1 Effect of number of samples
We observe that over the course of the evolutionary algorithm, with the parameter settings $B=10,IT=400,E=10$, a total of $B\times IT\times E=40,000$ different transit networks are considered. By comparison, LC-100 only considers 100 solutions. It could be that NEA’s superiority to LC-100 is only due to its considering more solutions. To test this, we ran LC-40k, in which we sample $40,000$ solutions from the learned constructor, and picks the lowest-cost of those. Comparing LC-100 and LC-40k in Figure 3, we see that across all three cities and all values of $\alpha$, LC-40k performs very similarly to LC-100, improving on it only slightly in comparison with the improvement given by EA or NEA. From this we conclude that the number of solutions considered is not on its own an important factor in EA’s and NEA’s performance: much more important is the evolutionary algorithm that guides the search through solution space.
@@ 5.2.2 Contribution of type-2 mutator
We now consider the impact of the type-2 mutator on NEA. This operator is the same between EA and NEA and is not a learned function. To understand how important it is to NEA’s performance, we ran “all-1 NEA”, a variant of NEA in which only the neural mutator is used, and not the type-2 mutator. The results are shown in Figure 4, along with the same curves for LC-100 and NEA as in Figure 3, for comparison. It is clear that NEA and all-1 NEA perform very similarly in most scenarios. The differences are most notable at the extreme values of $\alpha$ (at either end of each curve), where we see that all-1 NEA underperforms NEA, achieving higher values of the relevant cost function component; but at intermediate $\alpha$ values, NEA and all-1 NEA’s trip-time-versus-operating-cost curves overlap. It appears that the minor adjustments to existing routes made by the type-2 mutator are more important at extremes of $\alpha$, where the neural net policy has been pushed to the extremes of its behaviour.
While the type-2 mutator’s contribution is small, it does help performance at extreme values and does not appear to worsen it at intermediate values, so it is worth retaining as a component of the algorithm.
@@ 5.2.3 Importance of learned heuristics
We observe that the type-1 mutator used in EA and the neural mutator used in NEA differ in the space of changes each is capable of making. The type-1 mutator can only add shortest paths as routes, while the neural mutator may compose multiple shortest paths into its new routes. This structural difference, and not the quality of heuristics learned by the policy $\pi_{\theta}$, may be what gives NEA its advantage over EA.
To test this conjecture, we ran a variant of NEA in which the learned policy is replaced by a policy that chooses actions uniformly at random: $\pi_{rand}(a|s_{t})=\frac{1}{|\mathcal{A}_{t}|}\;\forall\;a\in\mathcal{A}_{t}$. We call this the Path-combining EA (PC-EA). Since we wanted to gauge the performance of this variant in isolation from the learned policy, we did not use LC-100 with a learned policy to generate the initial solution $\mathcal{R}_{0}$. Instead, we used this random policy in LC-100 to generate $\mathcal{R}_{0}$ for PC-EA.
The results are shown in Figure 5, along with the same LC-100, EA, and NEA curves as in Figure 3. We see that for most values of $\alpha$, PC-EA performs less well than both EA and NEA. On the other hand, at the extreme of $\alpha=0.1$ it performs better than both NEA and EA, decreasing average trip time $C_{p}$ to about two minutes less than NEA on average. PC-EA’s poor performance at low $\alpha$ makes sense: its ability to replace routes with composites of shortest paths biases it towards making longer routes, moreso than EA. This is an advantage at $\alpha=1.0$ but a disadvantage at $\alpha=0.0$ (and, as we observe here, at intermediate $\alpha$). It is interesting, though, that PC-EA outperforms NEA at $\alpha=1.0$, as these two share the same space of possible actions. We would expect that if this action space enables the degree of performance shown by PC-EA, then the neural net policy used in NEA should be able to learn to reach that performance as well. Yet somehow in this scenario it fails to do so. Nonetheless, NEA performs better overall than PC-EA, and so we leave the exploration of this mystery to future work. For now we merely note that composing shortest paths appears to be a very useful heuristic for route construction in its own right, even when done purely randomly.
@@ 5.3 Comparisons with prior work
During the experiments of subsection 5.1, we observed that after 400 iterations NEA’s best-solution quality was still improving. In order to make a fairer comparison with other methods from the literature on the Mandl and Mumford benchmarks, we performed a further set of experiments where we ran NEA with $IT=4,000$ instead of $400$, for $\alpha=0.0$ and $\alpha=1.0$. In addition, because we observed better performance from PC-EA for $\alpha$ near $1.0$, we also ran PC-EA with $IT=4,000$ for $\alpha=1.0$, and we include these results as well.
@@
Table 3 and Table 4 present the results of these experiments on Mandl and the four Mumford cities, alongside results reported on these cities in comparable recent work. All results we report for NEA and PC-EA are averaged over ten runs with different random seeds; the same set of ten trained models used in the experiments of subsection 5.1 and subsection 5.2 are used here in NEA. Results for other methods are as reported by their authors. In addition to average trip time $C_{p}$ and total route time $C_{o}$, the tables present $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$. For $i\in[0,2]$, $d_{i}$ is the percentage of all passenger trips that require $i$ transfers between transit routes; while $d_{un}$ is the percentage of trips that require more than 2 transfers ($d_{un}=100-(d_{0}+d_{1}+d_{2})$).
@@
In these comparisons, we do not include results from Ahmed et al. (2019). While they report impressive results on these benchmarks, they do not provide the values used for the two parameters to their algorithm. In our earlier work Holliday and Dudek (2023), theirs was one of several methods we implemented in order to test our proposed initialization scheme. We discovered parameter values on our own that gave results comparable to other work up to 2019, but unfortunately after much effort and correspondence with one of the authors of Ahmed et al. (2019), we were unable to replicate their reported results. For this reason, we do not include their reported results here.
@@
On a desktop computer with a 2.4 GHz Intel i9-12900F processor and an NVIDIA RTX 3090 graphics processing unit (used to accelerate our neural net computations), NEA takes about 3 hours for each run on Mumford3, the largest environment, while the PC-EA runs take about 2.5 hours. By comparison, John et al. (2014) and Hüsselmann et al. (2023)’s NSGA-II variant, which each employ a genetic algorithm with a population of 200 solutions, take more than two days to run on Mumford3. Kılıç and Gök (2014) report that their procedure takes eight hours just to construct the initial solution for Mumford3, and don’t report the running time for the subsequent optimization. Hüsselmann et al. (2023)’s DBMOSA variant, meanwhile, takes 7 hours and 52 minutes to run on Mumford3.
@@
These reported running times are not directly comparable as the experiments were not run on identical hardware, but the differences are broadly indicative of the differences in speed of these methods. The running times of these methods is mainly a function of the metaheuristic algorithm used, rather than the mutation operators that alter solutions at each step. Genetic algorithms like NSGA-II, with large populations as used in some of these methods, are known to be very time-consuming because of the large number of solutions they must modify and evaluate at each step. But their search of solution space is correspondingly more exhaustive than single-solution methods such as simulated annealing, or an evolutionary algorithm with a small population ($B=10$) as we use in our own experiments. It is therefore to be expected that they would achieve lower final costs in exchange for their greater run-time. We are interested here in the quality of the low-level heuristics we propose, rather than the metaheuristic algorithm, and this must be kept in mind as we discuss the results of this section.
@@
Table 3: Passenger-perspective results. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the number of trips satisfied with number of transfers $i$, while $d_{un}$ is the number of trips satsified with 3 or more transfers. Arrows next to each quantity indicate which of increase or decrease is desirable. Bolded values in $C_{p}$, $d_{0}$, and $d_{un}$ columns are the best on that environment.
Table 3 shows the passenger perspective results. On each city except Mumford1, the NSGA-II variant of Hüsselmann et al. (2023) has the best final performance; it is also among the most time-consuming of the algorithms reported on here. PC-EA and NEA both perform poorly on the smallest two cities, Mandl and Mumford0, but their relative performance improves as the size of the city increases. On both Mumford2 and Mumford3, PC-EA’s performance is very close to that of Hüsselmann et al. (2023)’s NSGA-II, and better than all other methods listed, despite the relatively under-powered metaheuristic that drives it. In particular, it outperforms Hüsselmann et al. (2023)’s DBMOSA, which uses the same low-level heuristics as the NSGA-II with a faster but less-exhaustive metaheuristic. That PC-EA exceeds its performance on Mumford2 and Mumford3 is evidence that the low-level heuristics used in PC-EA are better for the passenger perspective than those proposed by Hüsselmann et al. (2023). This is especially the case given that DBMOSA is still a more powerful metaheuristic, making use of a hyper-heuristic to adapt the probability that each of its low-level heuristics will be applied over the run. By contrast, PC-EA always applies both of its heuristics equally often.
NEA does not perform as well as PC-EA on (aligning with what we observed in subsection 5.2), but still shows good performance on Mumford2 and 3, here outperforming all of these methods that were published prior to 2022.
Table 4 makes the same comparison for the operator perspective. Kılıç and Gök (2014) does not report results for the operator perspective and as such we do not include it in this table. Similarly to the passenger-perspective results, our methods underperform on the smallest cities (Mandl and Mumford0) but perform well on larger ones. Strikingly, despite NEA’s under-powered metaheuristic, it achieves the lowest value of total route time $C_{o}$ out of all methods on Mumford1, 2, and 3, improving even on Hüsselmann et al. (2023)’s NSGA-II with its much higher computational cost. Evidently, the learned policy functions very well as a low-level heuristic at low values of $\alpha$, where the premium is on keeping routes short, though less well when the premium is on shortening passenger trips.

### v2

@@ 5 MUMFORD EXPERIMENTS
We first evaluated our method on the Mandl (1980) and Mumford (2013b) city datasets, two popular benchmarks for evaluating NDP algorithms (Mumford, 2013a; John et al., 2014; Kılıç and Gök, 2014; Ahmed et al., 2019). The Mandl dataset is one small synthetic city, while the Mumford dataset consists of four synthetic cities, labelled Mumford0 through Mumford3, and gives values of $S$, $MIN$, and $MAX$ to use when benchmarking on each city. The values $n$, $S$, $MIN$, and $MAX$ for Mumford1, Mumford2, and Mumford3 are taken from three different real-world cities and their existing transit networks, giving the dataset a degree of realism. Details of these benchmarks, and the parameters we use when evaluating on them, are given in Table 1.
@@
City # nodes $n$ street edges $|\mathcal{E}_{s}|$ routes $S$ $MIN$ $MAX$ Area (km2)
@@ 5.1 Comparison with Baseline Evolutionary Algorithm
We run each algorithm under consideration on all five of these synthetic cities over eleven different values of $\alpha$, ranging from $0.0$ to $1.0$ in increments of $0.1$. This lets us observe how well the different methods perform under a range of possible preferences, from the extremes of the operator perspective ($\alpha=0.0$, caring only about $C_{o}$) and passenger perspective ($\alpha=1.0$, caring only about $C_{p}$) to a range of intermediate perspectives. The constraint weight is set as $\beta=5.0$ in all experiments, the same value used in training the policies. We found that this was sufficient to prevent any of the constraints in section 3 from being violated by any transit network produced in our experiments.
Because these algorithms are stochastic, for each city we perform ten runs of each algorithm with ten different random seeds. For algorithms that make use of a learned policy, ten separate policies were trained with the same set of random seeds (but using the same training dataset), and each was used when running algorithms with the corresponding random seed. The values reported are statistics computed over the ten runs.
We first compared our neural evolutionary algorithm (NEA) with our baseline evolutionary algorithm (EA). We also compare both with the initial networks from LC-100, to see how much improvement each algorithm makes over the initial networks. In the EA and NEA runs, the same parameter settings of $B=10,IT=400,E=10$ were used, following the values used in Nikolić and Teodorović (2013).
Table 2 displays the mean and standard deviation of the cost $C(\mathcal{C},\mathcal{R})$ achieved by each algorithm on the Mandl and Mumford benchmarks for the operator perspective ($\alpha=0.0$), the passenger perspective ($\alpha=1.0$), and a balanced perspective ($\alpha=0.5$). We see that on Mandl, the smallest of the five cities, EA and NEA perform virtually identically for the operator and passenger perspectives, while NEA offers a slight improvement for the balanced perspective. But for the larger cities of the Mumford dataset, NEA performs considerably better than EA for both the operator and balanced perspectives. On Mumford3, the largest of the five cities, NEA solutions have 14% lower average cost for the operator perspective and 6% lower for the balanced perspective than EA.
For the passenger perspective NEA’s advantage over EA is smaller, but it still outperforms EA on Mumford1 and Mumford3 and performs comparably on Mumford2. NEA’s solutions have 1% lower average cost on Mumford3 than EA’s, but the average costs of NEA and EA are within one standard deviation of each other.
@@
Figure 3 displays the average trip time $C_{p}$ and total route time $C_{o}$ achieved by each algorithm on the three largest Mumford cities, the three which are each based on a real city’s statistics, over eleven $\alpha$ values evenly spaced over the range $[0,1]$ in increments of 0.1. There is a necessary trade-off between $C_{p}$ and $C_{o}$, and as we would expect, as $\alpha$ increases, $C_{o}$ increases and $C_{p}$ decreases for each algorithm’s output. We also observe that for intermediate values of $\alpha$, NEA pushes $C_{o}$ much lower than LC-100, at the cost of increases in $C_{p}$. This behaviour is desirable, because as the figure shows, NEA achieves an overall larger range of outcomes than LC-100 - for any network $\mathcal{R}$ from LC-100, there is some value of $\alpha$ for which NEA produces a network $\mathcal{R}^{\prime}$ which dominates $\mathcal{R}$. Figure 3 also shows results for an additional algorithm, LC-40k, which we discuss in subsection 5.2.
On all three of Mumford1 through Mumford3, we observe a common pattern: EA and NEA perform very similarly at $\alpha=1.0$ (the leftmost point on each curve), but a significant performance gap forms as $\alpha$ decreases - consistent with what we see in Table 2. We also observe that LC-100 favours reducing $C_{p}$ over $C_{o}$, with its points clustered higher and more leftwards than most of the points with corresponding $\alpha$ values on the other curves. Both EA and NEA achieve only very small decreases in $C_{p}$ on LC-100’s initial solutions at $\alpha=1.0$, and they do so by increasing $C_{o}$ considerably.
@@
Figure 3: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis), across values of $\alpha$ over the range $[0,1]$ in increments of $0.1$, from $\alpha=0.0$ at lower-right to $1.0$ at upper-left, for transit networks from LC-100, LC-40k, EA, and NEA. Each point shows the mean $C_{p}$ and $C_{o}$ over 10 random seeds for one value of $\alpha$, and bars around each point indicate one standard deviation on each axis. Lines linking pairs of points indicate that they represent consecutive $\alpha$ values. Lower values of $C_{o}$ and $C_{p}$ are better, so the down-and-leftward direction in each plot represents improvement.
@@
Figure 4: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis) achieved by all-1 NEA, plotted along with NEA (repeated from Figure 3) for comparison.
@@
Figure 5: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis) achieved by PC-EA, plotted along with LC-100, EA, and NEA (repeated from Figure 3) for comparison.
@@ 5.2 Ablation Studies
To better understand the contribution of various components of our method, we performed three sets of ablation studies. These were conducted over the three realistic Mumford cities (1, 2, and 3) and over eleven $\alpha$ values evenly spaced over the range $[0,1]$ in increments of 0.1.
@@ 5.2.1 Effect of number of samples
We note that over the course of the evolutionary algorithm, with our parameter settings $B=10,IT=400,E=10$, a total of $B\times IT\times E=40,000$ different transit networks are considered. By comparison, LC-100 only considers 100 networks. It could be that NEA’s superiority to LC-100 is only due to its considering more networks. To test this, we ran LC-40k, in which we sample $40,000$ networks from the learned-construction algorithm, and pick the lowest-cost network. Comparing LC-100 and LC-40k in Figure 3, we see that across all three cities and all values of $\alpha$, LC-40k performs very similarly to LC-100, improving on it only slightly in comparison with the larger improvements given by EA or NEA. From this we conclude that the number of networks considered is not on its own an important factor in EA’s and NEA’s performance: much more important is the evolutionary algorithm that guides the search of possible networks.
@@ 5.2.2 Contribution of type-2 mutator
We next considered the impact of the type-2 mutator heuristic on NEA. This low-level heuristic is the same between EA and NEA and is not a learned function. To understand how important it is to NEA’s performance, we ran “all-1 NEA”, a variant of NEA in which only the neural mutator is used, and not the type-2 mutator. The results are shown in Figure 4, along with the curve for NEA as in Figure 3, for comparison. It is clear that NEA and all-1 NEA perform very similarly in most scenarios. The differences are most notable at the extreme values of $\alpha$ (at either end of each curve), where we see that all-1 NEA underperforms NEA, achieving higher values of the relevant cost function component; but at intermediate $\alpha$ values on Mumford1 and Mumford2, NEA and all-1 NEA’s trip-time-versus-operating-cost curves overlap. It appears that the minor adjustments to existing routes made by the type-2 mutator are more important at extremes of $\alpha$, where the neural net policy has been pushed to the extremes of its behaviour.
On Mumford3, we observe that the curves do not overlap as cleanly; particularly for $\alpha\leq 0.5$, NEA seems to slightly outperform all-1 NEA. So the contribution of the type-2 mutator appears to grow as the city gets larger and more challenging. This may be because as the city gets larger, the space of possible routes grows larger, so the neural mutator’s changes get more dramatic in comparison with the single-node adjustments of the type-2 mutator. The type-2 mutator’s class of changes becomes more distinct from those of the neural mutator, possibly making it more important.
While the type-2 mutator’s contribution is small, it does help performance in some cases. We conclude that it is a useful component of the algorithm.
@@ 5.2.3 Importance of learned heuristics
We note that the type-1 mutator used in EA and the neural mutator used in NEA differ in the space of changes each is capable of making. The type-1 mutator can only add shortest paths as routes, while the neural mutator may compose multiple shortest paths into its new routes. It may be that this structural difference, as opposed to the quality of heuristics learned by the policy $\pi_{\theta}$, is part of NEA’s advantage over EA.
To test this conjecture, we ran a variant of NEA in which the learned policy $\pi_{\theta}$ is replaced by a policy $\pi_{\text{random}}$ that chooses actions uniformly at random:
$\pi_{\text{random}}(a|s_{t})=\frac{1}{|\mathcal{A}_{t}|}\;\forall\;a\in\mathcal{A}_{t}$
We call this variant the path-combining evolutionary algorithm (PC-EA). Since we wanted to gauge the performance of this variant in isolation from the learned policy, we did not use LC-100 with $\pi_{\theta}$ to generate the initial network $\mathcal{R}_{0}$. Instead, we used $\pi_{\text{random}}$ in LC-100 to generate $\mathcal{R}_{0}$ for PC-EA.
The results are shown in Figure 5, along with the same LC-100, EA, and NEA curves as in Figure 3. For most values of $\alpha$, PC-EA performs less well than both EA and NEA. But at the extreme of $\alpha=1.0$ it performs better than both NEA and EA, decreasing average trip time $C_{p}$ by between one and two minutes versus NEA on each city - an improvement of about 5% on Mumford1 and Mumford3 and about 3.5% on Mumford2.
PC-EA’s poor performance at low $\alpha$ makes sense: its ability to replace routes with composites of shortest paths biases it towards making longer routes, moreso than EA, and unlike $\pi_{\theta}$, $\pi_{\text{random}}$ cannot learn to prefer halting early in such cases. This is an advantage over EA at $\alpha=1.0$ but a disadvantage at $\alpha$ less than about $0.8$. It is interesting, though, that PC-EA outperforms NEA at $\alpha=1.0$, as these two share the same space of possible actions. We would expect that the neural net policy used in NEA should be able to learn to reach that performance of PC-EA simply by learning to choose actions uniformly at random. Yet somehow it fails to do so.
Nonetheless, NEA performs better overall than PC-EA, and so we leave the exploration of this mystery to future work. For now we merely note that composing shortest paths is a good heuristic for route construction at high $\alpha$, but is worse than simply choosing shortest paths at low $\alpha$; so the advantage of NEA comes primarily from what the policy $\pi_{\theta}$ learns during training.
@@ 5.3 Comparisons with prior work
While conducting the experiments of subsection 5.1, we observed that after 400 iterations the cost of NEA’s best-so-far network $\mathcal{R}_{\text{best}}$ was still decreasing from one iteration to the next. In order to make a fairer comparison with other methods from the literature on the Mandl and Mumford benchmarks, we performed a further set of experiments where we ran NEA with $IT=4,000$ instead of $400$, for the operator perspective ($\alpha=0.0$) and the passenger perspective ($\alpha=1.0$). In addition, because we observed better performance from PC-EA for $\alpha$ near $1.0$, we also ran PC-EA with $IT=4,000$ for $\alpha=1.0$, and we include these results in the comparison.
@@
Table 3 and Table 4 present the results of these experiments on the Mandl and Mumford benchmarks, alongside results reported on these benchmarks in comparable recent work. As elsewhere, the results we report for NEA and PC-EA are averaged over ten runs with different random seeds; the same set of ten trained policies used in the experiments of subsection 5.1 and subsection 5.2 are used here in NEA. Results from other work are as reported in that work. In addition to the average trip time $C_{p}$ and total route time $C_{o}$, the tables present metrics of how many transfers between routes were required by the networks. These are labelled $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$. $d_{i}$ with $i\in\{0,1,2\}$ is the percentage of all passenger trips that required $i$ transfers between routes; while $d_{un}$ is the percentage of trips that require more than 2 transfers ($d_{un}=100-(d_{0}+d_{1}+d_{2})$).
@@
In these comparisons, we do not include results from Ahmed et al. (2019). While they reported results that set a new state of the art on these benchmarks, they do not provide the values used for the two parameters to their algorithm. In our earlier work Holliday and Dudek (2023), theirs was one of several methods we implemented in order to test our proposed initialization scheme. We discovered parameter values on our own that gave results comparable to other work up to 2019. But unfortunately, after much effort and correspondence with one of the authors of that work, we were unable to replicate their new state-of-the-art results. For this reason, we do not include their reported results here.
@@
Before discussing these results, a brief word should be said about computation time. On a desktop computer with a 2.4 GHz Intel i9-12900F processor and an NVIDIA RTX 3090 graphics processing unit (used to accelerate our neural net computations), NEA takes about 3 hours for each run on Mumford3, the largest environment, while the PC-EA runs take about 2.5 hours. By comparison, John et al. (2014) and Hüsselmann et al. (2023) both use a variant of NSGA-II, a genetic algorithm, with a population of 200 networks, which in both cases takes more than two days to run on Mumford3. Hüsselmann et al. (2023)’s DBMOSA variant, meanwhile, takes 7 hours and 52 minutes to run on Mumford3. Kılıç and Gök (2014) report that their procedure takes eight hours just to construct the initial network for Mumford3, and don’t report the running time for the subsequent optimization.
@@
These reported running times are not directly comparable as the experiments were not run on identical hardware, but the differences are broadly indicative of the differences in speed of these methods. The running times of these methods is mainly a function of the metaheuristic algorithm used, rather than the low-level heuristics used in the algorithm. Genetic algorithms like NSGA-II, with large populations as used in some of these methods, are very time-consuming because of the large number of networks they must modify and evaluate at each step. But their search of solution space is correspondingly more exhaustive than single-solution methods such as simulated annealing, or an evolutionary algorithm with a small population ($B=10$) as we use in our own experiments. It is therefore to be expected that they would achieve lower final costs in exchange for their greater run-time. We are interested here in the quality of the low-level heuristics we propose, rather than the metaheuristic algorithm, and this must be kept in mind as we discuss the results of this section.
@@
Table 3: Passenger-perspective results. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the percentage of trips satisfied with number of transfers $i$, while $d_{un}$ is the percentage of trips satsified with 3 or more transfers. Arrows next to each quantity indicate which of increase or decrease is desirable. Bolded values in $C_{p}$, $d_{0}$, and $d_{un}$ columns are the best on that environment.
5.3.1 Passenger-perspective results
Table 3 shows the passenger perspective results alongside results from other work. On each city except Mumford1, the NSGA-II variant of Hüsselmann et al. (2023) has the best final performance; it is also among the most time-consuming of the algorithms reported on here. PC-EA and NEA both perform poorly on the smallest two cities, Mandl and Mumford0, but their relative performance improves as the size of the city increases. On both Mumford2 and Mumford3, PC-EA’s performance is very close to that of Hüsselmann et al. (2023)’s NSGA-II, and better than all other methods listed, despite the relatively under-powered metaheuristic that drives it.
PC-EA also outperforms Hüsselmann et al. (2023)’s DBMOSA, which uses the same low-level heuristics as their NSGA-II with a faster but less-exhaustive metaheuristic. That PC-EA exceeds DBMOSA’s performance on Mumford2 and Mumford3 is evidence that the low-level heuristics used in PC-EA are better for the passenger perspective than those proposed by Hüsselmann et al. (2023). This is especially the case given that DBMOSA is still a more sophisticated metaheuristic algorithm than ours, with ten different low-level heuristics and a hyper-heuristic that adapts the rate at which each low-level heuristic is applied over the run.
NEA does not perform as well as PC-EA on the passenger perspective (aligning with what we observed in subsection 5.2), but still shows good performance on Mumford2 and 3, outperforming all of these methods that were published prior to 2022.
5.3.2 Operator-perspective results
Table 4 shows the operator-perspective results alongside results from other work. Kılıç and Gök (2014) do not report results for the operator perspective and as such we do not include their work in Table 4. Similarly to the passenger-perspective results, our methods underperform on the smallest cities (Mandl and Mumford0) but perform well on larger ones. Strikingly, NEA achieves the lowest value of total route time $C_{o}$ out of all methods on Mumford1, 2, and 3, improving even on Hüsselmann et al. (2023)’s NSGA-II. Evidently, the learned policy functions very well as a low-level heuristic at low values of $\alpha$, where the premium is on keeping routes short.